A market demand function is P= 100 - Q. MC = 40. Total revenues =P*Q = (100 - Q)*Q= 100Q – Q2. Therefore Marginal Revenue = dCan you please help me out with this problem? Thank you!!!
Answer #1

P=100- Q and dre=yo price equation At Р= мс 100-Q = 4o Q = 100-40 P=60 units. Aituting a into P = loorbo P =$40 TR= p.Q = (10

100-20 = 40 20 = 100-40 Q = 60 2 c Q = 30 units. Substituting a into price eqreation p = loor30 P = $70 Cournot model: From I

Firma (BR2) The= Pla, = [100-Q- Qe Q2 TR₂ = 100 - QQ - Q² STR 2 clooll) - a (0) - QQ2 MR₂ = 100 - Q - 20 profit-maximizing co

Q = 20 units. substituting az into BR, 20, +20=60 Q =60-20 2 a = un Q = 20 units. substituting Q, and a 2 into p equiti O equ